OpenStudy (anonymous):

Two square pyramids have the same volume. The first pyramid has base edge 20 in. and height 21 in. The second pyramid has height 84 in. What is the base edge of the second pyramid? 21 in. 42 in. 28 in. 10in

OpenStudy (abmon98):

Vof Pyramid=1/3*b^2*h V1=V2 1/3*(20)^2*21=1/3*b*84

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\)